# Metric Differential Privacy at the User-Level

### v1

@@ 6.2.2. Algorithms in the Central Model
For the metric space $\mathcal{X}=\mathcal{B}\times\mathcal{C}$, the Laplace mechanism described above satisfies $(\varepsilon,0)$-user level DP, and produces an estimate $\tilde{F}$ such that
$\max_{K}\mathbb{E}[d_{\textsf{EM}}(\tilde{K},\tilde{F})]\leq O\left(\tfrac{k}{n\varepsilon}\right).$
@@
It is possible to adapt FreqEstLocal to bounded central $d_{\textsf{EM}}$-DP by simply pretending to be one user who holds the global dataset $K_{G}$. The privacy analysis of Thm. 4.4, and the utility analysis of Thm. 6.6 may be combined for the following corollary:
@@
For the metric space $\mathcal{X}=\mathcal{B}\times\mathcal{C}$, FreqEstLocal with the mechanism $\mathcal{A}=\textsf{GKRR}_{\alpha_{0}}$ with $\alpha_{0}$ given in (7) satisfies $(\alpha,\delta)$-$d_{\textsf{EM}}$ DP in the central model and returns an estimator $\tilde{F}$ with error given in (7).
@@
$\max_{K}\mathbb{E}[d_{\textsf{EM}}(\tilde{F},\tilde{K})]\leq 4\tfrac{\sqrt{k^{3}}}{\sqrt{mn}}+64\tfrac{\sqrt{k^{3}}}{\alpha n}\sqrt{\ln(4m\exp(\alpha)/\delta)}.$
The same sampling error is present, but the cost of privacy is reduced from a $\frac{1}{\sqrt{n}}$ dependence in Thm. 6.6 to just $\frac{1}{n}$. To compare just the cost of privacy in Corollary 6.6 to Lemma 6.7, we will assume we are in the regime $n\leq\frac{m}{\alpha}$. Then, the cost in Corollary 6.8 is $O(\frac{\sqrt{k^{3}}}{\alpha n})\sqrt{\max\{\ln(\frac{m}{\delta}),\alpha\}}$. The error of Corollary 6.8 will be less when
@@
7. Related Work
@@
Item-level DP. DP was originally considered at the item-level (Dwork, 2006), where a privacy guarantee is made when one item in the sensitive dataset is changed. Of the most relevance to our setting are results in distribution estimation (Hay et al., 2009; Xu et al., 2013; Suresh, 2019); these results study more complex estimation problems than frequency. Also, we consider linear query release, for which there is a long line of work (Hardt and Talwar, 2010; Bhaskara et al., 2012; Nikolov et al., 2013; Blum et al., 2013; Li et al., 2015). The mechanism in (Hardt and Talwar, 2010) is often optimal and easy to adapt to our setting; we compare our algorithms with it.
@@
User-level DP. With a vast increase data collected about users, user-level privacy is gaining more interest (Amin et al., 2019; Narayanan et al., 2022; Bassily and Sun, 2023; Levy et al., 2021; Liu et al., 2020; Cummings et al., 2022). The most relevant work to ours on user-level private mean estimation (Cummings et al., 2022) and histogram estimation (Liu et al., 2023; Acharya et al., 2023), though these problems are again more complex than the ones we study. Another related area is the problem of deciding the amount of data to pick from each user in cases where the users have different amounts of data (Amin et al., 2019; Liu et al., 2023; Cummings et al., 2022), which is related to our unbounded DP setup. These techniques apply to more specialized settings than our general blackbox reduction, and they are not immediately comparable.
@@
Local DP. Local DP has also received lots of attention recently. The most results to our work is locally-private linear query release  (Duchi et al., 2013; Bassily, 2019) and distribution estimation (Duchi et al., 2013; Kairouz et al., 2016; Acharya et al., 2018; Chen et al., 2020; Acharya et al., 2023). We directly compare our work to the optimal algorithms in (Bassily, 2019) and (Chen et al., 2020) for our problems, which can be adapted to user-level DP easily. The other related line of work is privacy amplification from the local model to the central model, given access to a trusted shuffler (Erlingsson et al., 2019; Girgis et al., 2021; Feldman et al., 2022). We extend the state-of-the-art analysis in (Feldman et al., 2022) to general metric DP.
@@
Metric DP. Metric DP was first proposed in (Chatzikokolakis et al., 2013) in the central model. In the local model, this has led to work on releasing numeric data (Roy Chowdhury et al., 2022), location data (Andrés et al., 2013; Bordenabe et al., 2014; Chatzikokolakis et al., 2015; Weggenmann and Kerschbaum, 2021) and text (Feyisetan et al., 2019, 2020; Feyisetan and Kasiviswanathan, 2021; Imola et al., 2022). Unlike these works, we consider privacy in a general metric space. The most related work is that of (Fernandes et al., 2019), which proposes metric DP based on the $d_{\textsf{EM}}$ for releasing text embeddings. As explained in the introduction, we consider a much more general setting than (Fernandes et al., 2019).
8. Conclusion
@@
We have proposed metric DP at the user level using the earth-mover’s distance $d_{\textsf{EM}}$. This captures both the magnitude and structural aspects of changes in the data, resulting in a tailored privacy semantic. We have designed two novel privacy mechanisms under $d_{\textsf{EM}}$-DP which improves the utility over standard DP. Additionally, we have shown that general (unbounded) $d_{\textsf{EM}}$-DP can be reduced to the simpler case (bounded) where all users have the same amount of data. Finally, we have demonstrated that $d_{\textsf{EM}}$-DP .
@@ Appendix B Omitted Proofs from Section 4
B.2. Proof of Lemma 4.2
See 4.2 In the local model, by Theorem 4.1, we have $\|q_{f}(K)-q_{f}(K^{\prime})\|\leq\ell$. By adding noise drawn from $\Gamma(d,\frac{1}{\alpha})$, it is known this satisfies $(\alpha,0)$-DP (Hardt and Talwar, 2010). In the bounded central setting, we have $\|q_{f}(K)-q_{f}(K^{\prime})\|\leq\frac{\ell}{n}$, and thus we may add noise drawn from $\Gamma(d,\frac{1}{n\alpha})$.
B.3. Proof of Theorem 4.3
@@ B.4.1. Preliminary Lemmas
(9) $\mathsf{Shuffle}(F_{1}(Y_{1}^{b}),F_{2}(Y_{2}),\ldots,F_{m}(Y_{m}))\\ =\mathcal{S}(\textbf{0}(Y_{1}^{b},\ldots,Y_{m}),\textbf{1}(Y_{1}^{b},\ldots,Y_{m})).$
@@
(10) $r(u)=Q(x_{1}^{0})\ \ \ \forall u\in U$
(11) $r(v)=Q(x_{1}^{1})\ \ \ \forall v\in V$
(12) $r(w)=Q(x_{\pi(w)})\ \ \ \forall w\in W.$
@@
Having reduced the shuffling problem to a divergence between two fixed probability distributions, we follow the method of (Feldman et al., 2022) to compute this divergence. We use the following two results:
@@
(Restatement of Lemma A.1 from (Feldman et al., 2022)): Suppose $p\geq\frac{16\ln(2/\delta)}{m}$, $C\sim Bin(m-1,p)$ and $A\sim Bin(C,\frac{1}{2})$. Define $P=(A+1,C-A)$ and $Q=(A,C-A+1)$. Then, $D_{e^{\varepsilon}}(P\|Q)\leq\delta$, where
@@
(Restatement of Theorem 2 from (Balle et al., 2018)) Let $P,Q$ be probability distributions satisfying $P=\nu M+(1-\nu)N$ and $Q=\nu M^{\prime}+(1-\nu)N$ for distributions $M,M^{\prime},N$ and $\nu\in[0,1]$. Given $\alpha\geq 1$, define $\alpha^{\prime}=1+\nu(\alpha-1)$ and $\beta=\frac{\alpha^{\prime}}{\alpha}$. Then,
@@
(Restatement of Theorem 2.5 from (Kairouz et al., 2015)) Let $P,Q$ be two distributions and $\alpha\geq 1$ be a parameter such that $D_{\alpha}(P\|Q)=0$. Then, there exist distributions $M,N$ such that
@@
B.4.2. Completing the proof of Lemma B.1
@@
(13) $\mathcal{A}(x_{1}^{0})=(1-\beta)Q(x_{1}^{0})+\beta Q(x_{1}^{1})$
(14) $\mathcal{A}(x_{1}^{1})=\beta Q(x_{1}^{0})+(1-\beta)Q(x_{1}^{1})$
@@
(15) $\mathcal{A}(x_{1}^{0})=(1-\gamma)R(x_{1}^{0},x_{j})+\gamma R^{\prime}(x_{1}^{0},x_{j})~{}~{}~{}\forall j\geq 2$
(16) $\mathcal{A}(x_{j})=\gamma R(x_{1}^{0},x_{j})+(1-\gamma)R^{\prime}(x_{1}^{0},x_{j})~{}~{}~{}\forall j\geq 2$
(17) $\mathcal{A}(x_{1}^{1})=(1-\gamma)R(x_{1}^{1},x_{j})+\gamma R^{\prime}(x_{1}^{1},x_{j})~{}~{}~{}\forall j\geq 2$
(18) $\mathcal{A}(x_{j})=\gamma R(x_{1}^{1},x_{j})+(1-\gamma)R^{\prime}(x_{1}^{1},x_{j})~{}~{}~{}\forall j\geq 2.$
where $\gamma=\frac{1}{1+\exp(\varepsilon_{0})}$. Subtracting 15 and 16, we obtain that
(19) $\mathcal{A}(x_{j})=\frac{\gamma}{1-\gamma}\mathcal{A}(x_{1}^{0})+\frac{1-2\gamma}{1-\gamma}R^{\prime}(x_{1}^{0},x_{j})~{}~{}~{}\forall j\geq 2,$
and likewise 17 and 18 imply
(20) $\mathcal{A}(x_{j})=\frac{\gamma}{1-\gamma}\mathcal{A}(x_{1}^{1})+\frac{1-2\gamma}{1-\gamma}R^{\prime}(x_{1}^{1},x_{j})~{}~{}~{}\forall j\geq 2.$
Taking the average of 19 and 20, we obtain
(21) $\mathcal{A}(x_{j})=\frac{\gamma}{2(1-\gamma)}\mathcal{A}(x_{1}^{0})+\frac{\gamma}{2(1-\gamma)}\mathcal{A}(x_{1}^{1})+\frac{1-2\gamma}{1-\gamma}Q(x_{j})~{}~{}~{}\forall j\geq 2,$
where $Q(x_{j})=\frac{1}{2}R^{\prime}(x_{1}^{0},x_{j})+\frac{1}{2}R^{\prime}(x_{1}^{1},x_{j})$. Now, equations 13 and 14 imply that
@@
(22) $\mathcal{A}(x_{j})=\frac{\gamma}{2(1-\gamma)}Q(x_{1}^{0})+\frac{\gamma}{2(1-\gamma)}Q(x_{1}^{1})+\frac{1-2\gamma}{1-\gamma}Q(x_{j})~{}~{}~{}\forall j\geq 2.$
@@
(23) $\tilde{L}=\mathrm{Shuffle}(\mathcal{A}(x_{1}),\ldots,\mathcal{A}(x_{m}))$
(24) $\tilde{L}^{\prime}=\mathrm{Shuffle}(\mathcal{A}(x_{\pi(i)}^{\prime}),\ldots,\mathcal{A}(x_{\pi(m)}^{\prime})).$

### v2

@@ 6.2.2 Algorithms in the Central Model
For the metric space $\mathcal{X}=\mathcal{B}\times\mathcal{C}$, the Laplace mechanism described above satisfies $(\varepsilon,0)$-user level DP, and produces an estimate $\tilde{H}$ such that
$\max_{K}\mathbb{E}[d_{\textsf{EM}}(\tilde{K},\tilde{H})]\leq O\left(\tfrac{k}{n\varepsilon}\right).$
@@
It is possible to adapt FreqEstLocal to bounded central $d_{\textsf{EM}}$-DP by simply pretending to be one user who holds the global dataset $K_{G}$. The privacy analysis of Theorem 4.4, and the utility analysis of Theorem 6.6 may be combined for the following corollary:
@@
For the metric space $\mathcal{X}=\mathcal{B}\times\mathcal{C}$, FreqEstLocal with $\mathcal{A}=\textsf{GKRR}_{\alpha_{0}}$ with $\alpha_{0}$ given in Eq. (7) satisfies $(\alpha,\delta)$-$d_{\textsf{EM}}$ DP in the central model and returns an estimator $\tilde{H}$ with error given in Eq. (9).
@@
$\max_{K}\mathbb{E}[d_{\textsf{EM}}(\tilde{H},\tilde{K})]\leq 4\tfrac{\sqrt{k^{3}}}{\sqrt{mn}}+64\tfrac{\sqrt{k^{3}}}{\alpha n}\sqrt{\ln(4m\exp(\alpha)/\delta)}.$
The same sampling error is present, but the cost of privacy is reduced from a $\frac{1}{\sqrt{n}}$ dependence in Theorem 6.6 to just $\frac{1}{n}$. To compare just the cost of privacy in Corollary 6.6 to Lemma 6.7, we will assume we are in the regime $n\leq\frac{m}{\alpha}$. Then, the cost in Corollary 6.8 is $O(\frac{\sqrt{k^{3}}}{\alpha n})\sqrt{\max\{\ln(\frac{m}{\delta}),\alpha\}}$. The error of Corollary 6.8 is less when
@@
Open Questions. For linear queries, one open question is whether it is possible to obtain error competitive with user-level DP under $(\alpha,0)$-$d_{\textsf{EM}}$-DP. For frequency estimation, an immediate open question is whether it is possible to reduce the super-linear dependence $k$ to one that matches that of Lemma 6.4, and whether an improvement in error can be made when $\alpha\leq\varepsilon^{2}k$.
New Proof Techniques. Frequency estimation is a classic problem that has been thoroughly studied in the standard DP literature. Despite the well-studied nature of the problem, the use of $d_{\textsf{EM}}$-DP enables a novel mechanism design. In particular, we can work with a transition matrix (corresponding to the private mechanism $\mathcal{A}$) that is less noisy. Consequently, we use a natural generalization of $k$-Randomized Response, which allows for better utility analysis than standard DP.
7 Related Work
@@
Item-level DP. DP was originally considered at the item-level (Dwork, 2006). Relevant to our setting are results in distribution estimation (Hay et al., 2009; Xu et al., 2013; Suresh, 2019); these results study more complex estimation problems than frequency. We also consider linear query release (Hardt and Talwar, 2010; Bhaskara et al., 2012; Nikolov et al., 2013; Blum et al., 2013; Li et al., 2015). The mechanism in Hardt and Talwar (2010) is often optimal and easy to adapt to our setting; we compare our algorithms with it.
@@
User-level DP. User-level privacy is gaining increasing interest (Amin et al., 2019; Narayanan et al., 2022; Bassily and Sun, 2023; Levy et al., 2021; Liu et al., 2020; Cummings et al., 2022). The most relevant work to ours involves user-level private mean estimation (Cummings et al., 2022) and histogram estimation (Liu et al., 2023; Acharya et al., 2023), though these problems are more complex than those we study. Another related area is deciding the amount of data to collect from each user when users have varying amounts of data (Amin et al., 2019; Liu et al., 2023; Cummings et al., 2022), which relates to our unbounded DP setup. These techniques apply to more specialized settings than our general blackbox reduction and are not immediately comparable.
@@
Local DP. The results most relevant to our work in local DP are locally-private linear query release  Duchi et al. (2013); Bassily (2019) and distribution estimation (Duchi et al., 2013; Kairouz et al., 2016; Acharya et al., 2018; Chen et al., 2020; Acharya et al., 2023). We directly compare our work to the optimal algorithms in Bassily (2019) and Chen et al. (2020) for our problems, which can be adapted to user-level DP easily. The other related line of work is privacy amplification via shuffling (Erlingsson et al., 2019; Girgis et al., 2021; Feldman et al., 2022). We extend the state-of-the-art analysis in Feldman et al. (2022) to general metric DP.
@@
Metric DP. Metric DP was first proposed in Chatzikokolakis et al. (2013) in the central model. In the local model, this has led to work on releasing numeric data (Roy Chowdhury et al., 2022), location data (Andrés et al., 2013; Bordenabe et al., 2014; Chatzikokolakis et al., 2015; Weggenmann and Kerschbaum, 2021) and text (Feyisetan et al., 2019, 2020; Feyisetan and Kasiviswanathan, 2021; Imola et al., 2022). Unlike these works, we consider privacy in a general metric space. The most related work is that of Fernandes et al. (2019), which proposes metric DP based on $d_{\textsf{EM}}$ for releasing text embeddings. As explained in the introduction, we consider a more general setting than Fernandes et al. (2019).
8 Interpretation of $d_{\textsf{EM}}$-DP
In this section, we elaborate on how to interpret the $d_{\textsf{EM}}$-DP guarantee (and metric DP in general) relative to standard DP. We start by discussing the advantages offered by $d_{\textsf{EM}}$-DP. As discussed in Section 1, the primary benefit of $d_{\textsf{EM}}$-DP is that it offers a more fine-grained and nuanced privacy definition compared to standard DP. This results in a more flexible privacy-utility trade-off that is better suited than standard DP for many real-world applications. In addition, $d_{\textsf{EM}}$-DP unlocks new proof techniques that may also be applicable to standard DP. Specifically, the $d_{\textsf{EM}}$ metric introduces couplings that need to be explicitly addressed in privacy analysis. For instance, in Section 5 we showed that sampling with replacement is a smooth projection by explicitly viewing sampling from two datasets in terms of a coupling between them. While standard DP privacy analysis often implicitly uses couplings, we believe that some of our proof techniques for explicitly handling general couplings could also be beneficial in the context of standard DP.
Next, let us understand the technical relation between metric DP and standard DP. Metric DP is essentially a relaxation of standard DP. Any mechanism that satisfies metric DP (user-level or item-level) also satisfies standard DP, albeit with a potentially higher privacy parameter. In particular, any mechanism satisfying $(\alpha,\delta)$-$d_{\mathcal{X}}$-DP also satisfies $(\alpha\cdot d_{max},\delta)$-DP, where $d_{max}$ is the maximum $d_{\mathcal{X}}$ distance between any two pairs of inputs. Conversely, any mechanism that satisfies $(\varepsilon,\delta)$-DP also satisfies $(\frac{\varepsilon}{d_{min}},\delta)$-$d_{\mathcal{X}}$-D,P where $d_{min}$ is the minimum $d_{\mathcal{X}}$ distance between any two pairs of inputs. Hence, although in theory one can translate between these two privacy guarantees, the translation is very loose. Tightly analyzing the privacy parameter under metric DP (whether $d_{\textsf{EM}}$ or otherwise) for an arbitrary mechanism that satisfies standard DP is non-trivial and there is no one-size-fits-all method to do so. For instance, Algorithm 1 can be instantiated via the Laplace and Gaussian mechanism – both classic standard DP mechanisms – under some conditions. However, as discussed in Section 4.1, a more complex analysis is required to evaluate privacy under $d_{\textsf{EM}}$-DP. Additionally, in terms of mechanism design, a mechanism optimized for $d_{\textsf{EM}}$-DP might not be ideal for standard DP and vice-versa. For instance, our proposed Algorithm 4 for performing frequency estimation may not work well under standard DP (i.e., have high privacy parameters).
In what follows, we outline three concrete scenarios, where a practitioner should prefer $d_{\textsf{EM}}$-DP over standard DP. First, if the practitioner has a prior on the sensitive data indicating that the distance between any two user’s data is overwhelmingly likely to be $<1$ (assuming all distances are normalized), then $d_{\textsf{EM}}$-DP is clearly the better choice. This is because such a prior makes the worst-case scenario of antipodal data pairs—where two users’ data are completely dissimilar (the case that standard DP safeguards against)—highly unlikely in practice. For instance, this scenario may arise in the context of location data when the data corresponds to location information of employees of the same firm. In this case, weekday locations will be the same across all users, leading to small pairwise $d_{\textsf{EM}}$ distances. Second, a practitioner should opt for $d_{\textsf{EM}}$-DP when it captures a more realistic privacy semantics of the underlying data. Although ideally, we would like to prevent any data leakage about an individual, this is unfortunately not feasible in practice due to the vast amount of auxiliary information already publicly available about every individual. For instance, most people’s occupations are publicly available on social media profiles. Returning to our location data example, data collected over a month would reveal routine patterns, such as a person’s workplace, which is already public and hence doesn’t require protection. Rather, the more sensitive information is short-term location data gathered over say the course of a day (which might reveal non-routine visit to a friend or hospital). In such a scenario, $d_{\textsf{EM}}$-DP would offer a better privacy-utility trade-off with more realistic privacy guarantees than standard DP. Third, $d_{\textsf{EM}}$-DP may be preferable if the practitioner is restricted to work within a low privacy parameter regime (for instance, due to some government guideline). This is because for the same privacy parameter (i.e., $\alpha=\varepsilon$), $d_{\textsf{EM}}$-DP can offer a stronger privacy guarantee than standard DP while maintaining the same utility for certain queries, such as linear queries (Section 4.1).
Finally, we conclude with some caveats regarding the use of metric DP. Metric DP assigns varying levels of sensitivity to different neighboring pairs. Specifically, smaller changes between neighboring pairs are considered more sensitive and are therefore protected with a higher privacy guarantee. However, this approach may not be suitable for all contexts. For instance, in the case of medical records, where the data between individuals can be vastly different, standard DP may offer better privacy protection. When adopting metric DP, it is crucial for practitioners to clearly define what is considered sensitive and what is not, and to engage in discussions about whether these definitions align with acceptable privacy semantics. This transparency enables users to make informed decisions about whether the privacy guarantees provided meet their needs.
9 Conclusion
@@
We have proposed metric DP at the user level using the earth-mover’s distance, $d_{\textsf{EM}}$. This captures both the magnitude and structural aspects of changes in the data, resulting in a tailored privacy semantic. We have designed two novel privacy mechanisms under $d_{\textsf{EM}}$-DP which improves the utility over standard DP. Additionally, we have shown that general (unbounded) $d_{\textsf{EM}}$-DP can be reduced to the simpler case (bounded) where all users have the same amount of data. Finally, we have demonstrated that $d_{\textsf{EM}}$-DP, when tailored to the application, can offer improved utility over standard DP.
@@ Appendix B Omitted Proofs from Section 4
B.2 Proof of Lemma 4.2
See 4.2 In the local model, by Theorem 4.1, we have $\|q_{f}(K)-q_{f}(K^{\prime})\|\leq\ell$. By adding noise drawn from $\Gamma(d,\frac{1}{\alpha})$, it is known this satisfies $(\alpha,0)$-DP Hardt and Talwar (2010). In the bounded central setting, we have $\|q_{f}(K)-q_{f}(K^{\prime})\|\leq\frac{\ell}{n}$, and thus we may add noise drawn from $\Gamma(d,\frac{1}{n\alpha})$.
B.3 Proof of Theorem 4.3
@@ B.4.1 Preliminary Lemmas
$\mathsf{Shuffle}(F_{1}(Y_{1}^{b}),F_{2}(Y_{2}),\ldots,F_{m}(Y_{m}))\\ =\mathcal{S}(\textbf{0}(Y_{1}^{b},\ldots,Y_{m}),\textbf{1}(Y_{1}^{b},\ldots,Y_{m})).$ (10)
@@
$r(u)=Q(x_{1}^{0})\ \ \ \forall u\in U$ (11)
$r(v)=Q(x_{1}^{1})\ \ \ \forall v\in V$ (12)
$r(w)=Q(x_{\pi(w)})\ \ \ \forall w\in W.$ (13)
@@
Having reduced the shuffling problem to a divergence between two fixed probability distributions, we follow the method of Feldman et al. (2022) to compute this divergence. We use the following two results:
@@
(Restatement of Lemma A.1 from Feldman et al. (2022)): Suppose $p\geq\frac{16\ln(2/\delta)}{m}$, $C\sim Bin(m-1,p)$ and $A\sim Bin(C,\frac{1}{2})$. Define $P=(A+1,C-A)$ and $Q=(A,C-A+1)$. Then, $D_{e^{\varepsilon}}(P\|Q)\leq\delta$, where
@@
(Restatement of Theorem 2 from Balle et al. (2018)) Let $P,Q$ be probability distributions satisfying $P=\nu M+(1-\nu)N$ and $Q=\nu M^{\prime}+(1-\nu)N$ for distributions $M,M^{\prime},N$ and $\nu\in[0,1]$. Given $\alpha\geq 1$, define $\alpha^{\prime}=1+\nu(\alpha-1)$ and $\beta=\frac{\alpha^{\prime}}{\alpha}$. Then,
@@
(Restatement of Theorem 2.5 from Kairouz et al. (2015)) Let $P,Q$ be two distributions and $\alpha\geq 1$ be a parameter such that $D_{\alpha}(P\|Q)=0$. Then, there exist distributions $M,N$ such that
@@
B.4.2 Completing the proof of Lemma B.1
@@
$\mathcal{A}(x_{1}^{0})=(1-\beta)Q(x_{1}^{0})+\beta Q(x_{1}^{1})$ (14)
$\mathcal{A}(x_{1}^{1})=\beta Q(x_{1}^{0})+(1-\beta)Q(x_{1}^{1})$ (15)
@@
$\mathcal{A}(x_{1}^{0})=(1-\gamma)R(x_{1}^{0},x_{j})+\gamma R^{\prime}(x_{1}^{0},x_{j})~{}~{}~{}\forall j\geq 2$ (16)
$\mathcal{A}(x_{j})=\gamma R(x_{1}^{0},x_{j})+(1-\gamma)R^{\prime}(x_{1}^{0},x_{j})~{}~{}~{}\forall j\geq 2$ (17)
$\mathcal{A}(x_{1}^{1})=(1-\gamma)R(x_{1}^{1},x_{j})+\gamma R^{\prime}(x_{1}^{1},x_{j})~{}~{}~{}\forall j\geq 2$ (18)
$\mathcal{A}(x_{j})=\gamma R(x_{1}^{1},x_{j})+(1-\gamma)R^{\prime}(x_{1}^{1},x_{j})~{}~{}~{}\forall j\geq 2.$ (19)
where $\gamma=\frac{1}{1+\exp(\varepsilon_{0})}$. Subtracting 16 and 17, we obtain that
$\mathcal{A}(x_{j})=\frac{\gamma}{1-\gamma}\mathcal{A}(x_{1}^{0})+\frac{1-2\gamma}{1-\gamma}R^{\prime}(x_{1}^{0},x_{j})~{}~{}~{}\forall j\geq 2,$ (20)
and likewise 18 and 19 imply
$\mathcal{A}(x_{j})=\frac{\gamma}{1-\gamma}\mathcal{A}(x_{1}^{1})+\frac{1-2\gamma}{1-\gamma}R^{\prime}(x_{1}^{1},x_{j})~{}~{}~{}\forall j\geq 2.$ (21)
Taking the average of 20 and 21, we obtain
$\mathcal{A}(x_{j})=\frac{\gamma}{2(1-\gamma)}\mathcal{A}(x_{1}^{0})+\frac{\gamma}{2(1-\gamma)}\mathcal{A}(x_{1}^{1})+\frac{1-2\gamma}{1-\gamma}Q(x_{j})~{}~{}~{}\forall j\geq 2,$ (22)
where $Q(x_{j})=\frac{1}{2}R^{\prime}(x_{1}^{0},x_{j})+\frac{1}{2}R^{\prime}(x_{1}^{1},x_{j})$. Now, equations 14 and 15 imply that
@@
$\mathcal{A}(x_{j})=\frac{\gamma}{2(1-\gamma)}Q(x_{1}^{0})+\frac{\gamma}{2(1-\gamma)}Q(x_{1}^{1})+\frac{1-2\gamma}{1-\gamma}Q(x_{j})~{}~{}~{}\forall j\geq 2.$ (23)
@@
$\tilde{L}=\mathrm{Shuffle}(\mathcal{A}(x_{1}),\ldots,\mathcal{A}(x_{m}))$ (24)
$\tilde{L}^{\prime}=\mathrm{Shuffle}(\mathcal{A}(x_{\pi(i)}^{\prime}),\ldots,\mathcal{A}(x_{\pi(m)}^{\prime})).$ (25)